OpenStudy (anonymous):

What is the inverse of f(x)= 2 -x?

OpenStudy (anonymous):

2+x ??

OpenStudy (anonymous):

f(-x) is what you are asking for right?

OpenStudy (anonymous):

o wait nm

OpenStudy (anonymous):

y = 2 -x 2 - y = x \[f^{-1}= 2 -x\]

OpenStudy (anonymous):

sorry its, f(x)= 2-x when f^-1..

OpenStudy (anonymous):

it is its own inverse

OpenStudy (anonymous):

\[y=2-x\iff x =2-y\]

OpenStudy (anonymous):

Thanks a lot Ishaan94, satellite73, and ksmith!